Larrys Building Supplies (LBS) is a local hardware store. LBS uses a perpetual inventory system. The following transactions2. Compute the gross profit percentage. (Round your answer to 1 decimal place.) Gross Profit PercentageRecord the return by a customer of unsatisfactory merchandise that was in perfect condition. A cash refund of $3,000 was giveRecord the collection of half of the balance of $2,500 owed by the customer within the discount period. Note: Enter debits beRecord the allowance of $950 granted to the customer. Note: Enter debits before credits. General Journal Transaction Debit Cr4. LBS is considering a contract to sell building supplies to a local home builder for $20,000. These materials will cost LBS
Answer #1
1] Sales = 500000+5000 = $        5,05,000
-Returns and allowances = 3000+950 = $             3,950
Net sales $        5,01,050
COGS = 224350-1900 = $        2,22,450
Gross profit = 501050-22450 = $        2,78,600
2] Gross profit percentage = Gross profit/Net sales = 278600/501050 = 55.6%
3] Sales $              3,000
Cash $            3,000
[To record payment of cash for return of goods]
Merchandise inventory $              1,900
COGS $            1,900
[To record cost of goods returned]
Cash $              2,500
Accounts receivable $            2,500
[To record cash received]
Note:
No discount rate is specified; only n/30 is mentioned.
Sales returns and allowances $                  950
Accounts receivable $                950
[To record returns and allowances]
4] Net sales = 501050+20000 = $        5,21,050
COGS = 222450+16000 = $        2,38,450
=Gross profit $        2,82,600
Gross profit percentage = 282600/521050 = 54.24%